# Real-world models for multiple term structures: a unifying HJM semimartingale framework

### v2

@@ 4.4. Existence of affine realizations
Since $\widehat{X}_{\cdot\wedge 1}=\widehat{X}^{1}\in\mathcal{X}_{1,[0,1]}$ and $\mathcal{X}_{1,[0,n]}\subseteq\mathcal{X}_{1}$, for every $n\in\mathbb{N}$, we have that $\widehat{X}_{\cdot\wedge 1}\in\mathcal{X}_{1}$. Proceeding by induction, suppose that $\widehat{X}_{\cdot\wedge n}\in\mathcal{X}_{1}$, for some $n\in\mathbb{N}$. Then, the definition of $\widehat{X}$ implies that
@@
$=\mathbf{1}_{\{t<n\}}(1+\widehat{X}_{t\wedge n})+\mathbf{1}_{\{t\geq n\}}\frac{1+\widehat{X}_{n}}{1+\widehat{X}^{n+1}_{n}}(1+\widehat{X}^{n+1}_{t}).$
Since $\widehat{X}^{n+1}\in\mathcal{X}_{1,[0,n+1]}\subseteq\mathcal{X}_{1}$ and the set $\mathcal{X}_{1}$ is fork-convex, it follows that $\widehat{X}_{\cdot\wedge(n+1)}\in\mathcal{X}_{1}$. We have thus shown that $\widehat{X}_{\cdot\wedge n}\in\mathcal{X}_{1}$, for all $n\in\mathbb{N}$. Since $\widehat{X}$ is a semimartingale and $\widehat{X}_{\cdot\wedge n}=\mathbf{1}_{[\kern-1.04996pt[0,n]\kern-1.04996pt]}\cdot\widehat{X}$, for all $n\in\mathbb{N}$, the closedness of $\mathcal{X}_{1}$ in the semimartingale topology implies that $\widehat{X}\in\mathcal{X}_{1}$.
Conversely, if a process $Z$ is a supermartingale deflator and $X\in\mathcal{X}_{1}$, then the supermartingale property of $Z(1+X)$ together with the fact that $Z_{0}\leq 1$ implies that $\mathbb{E}[Z_{T}(1+X_{T})]\leq 1$, for every $T\in\mathbb{R}_{+}$. This implies that the set $Z_{T}\mathcal{X}_{1}(T)$ is bounded in probability and, hence, NUPBR holds.
Finally, suppose that $Z$ is a local martingale deflator and let $n\in\mathbb{N}$ and $A\in\mathcal{A}^{n}$. By definition, $ZX^{A}$ is an $\mathbb{R}^{n}$-valued local martingale. For every $H\in L(X^{A})$ with $X:=H\cdot X^{A}\geq-1$, the process $ZX$ is a local martingale (see, e.g., [28, Lemma 4.2]). Since every non-negative local martingale with integrable initial value is a supermartingale, this implies that $Z(1+X)$ is a supermartingale, meaning that $Z$ is a supermartingale deflator for the set $\cup_{n\geq 1}\mathcal{X}^{n}_{1}$. By Fatou’s lemma, the same property holds for the closure $\mathcal{X}_{1}$ in the semimartingale topology, thus proving that $Z$ is a supermartingale deflator. ∎
@@ Appendix B Conditional expectation with respect to a Doléans measure
We recall the notion of conditional expectation with respect to a Doléans measure, as introduced in [34] (see also [35, Section III.3c]). Let $\mu$ be an integer-valued random measure on +×$\mathbb{R}_{+}\times E$ with compensator $\nu$. The positive Doléans measure $M_{\mu}$ on $(\Omega\times\mathbb{R}_{+}\times E,\mathcal{F}\otimes\mathcal{B}(\mathbb{R}_{+})\otimes\mathcal{B}(E))$ is defined by
@@
Suppose there exists a sequence $\{\tau_{n}\}_{n\in\mathbb{N}}$ of stopping times increasing a.s. to infinity as $n\to+\infty$ such that $\mathbb{E}[(|\varphi|\ast\mu)_{\tau_{n}}]<+\infty$, for all $n\in\mathbb{N}$. For each $n\in\mathbb{N}$, the function $H_{n}:\Omega\times\mathbb{R}_{+}\times E\to\{0,1\}$ defined by $H_{n}:=\mathbf{1}_{[\kern-1.04996pt[0,\tau_{n}]\kern-1.04996pt]\times E}$ is $\widetilde{\mathcal{P}}$-measurable and bounded. Therefore, property (B.1) implies that
@@
for all $n\in\mathbb{N}$, where in the last equality we made use of [35, Theorem II.1.8]. This shows that the process $M_{\mu}[|\varphi||\widetilde{\mathcal{P}}]\ast\nu$ is locally integrable. The converse implication can be shown in the same way. To prove the second part of the lemma, by localization we can assume that $\mathbb{E}[(|\varphi|\ast\mu)_{\infty}]<+\infty$. Let $\tau$ be an arbitrary stopping time. Similarly as above, letting the function $H$ be defined by $H:=\mathbf{1}_{[\kern-1.04996pt[0,\tau]\kern-1.04996pt]\times E}$,
@@
thus implying that the process $\varphi\ast\mu-M_{\mu}[\varphi|\widetilde{\mathcal{P}}]\ast\nu$ is a martingale. In view of [35, Theorem I.3.18], this suffices to deduce that $M_{\mu}[\varphi|\widetilde{\mathcal{P}}]\ast\nu$ is the compensator of $\varphi\ast\mu$. ∎
@@ Appendix D Properties of multi-dimensional Filipović spaces
    the translation semigroup $(S_{t})_{t\geq 0}$ is a $C_{0}$-semigroup on $H_{\rho}$;
@@
[48, Lemma 4.2] The pair $(H_{\rho},m)$, where $m:H_{\rho}\times H_{\rho}\to H_{\rho}$ denotes the pointwise multiplication $m(h,g):=hg=h\cdot g$, is a commutative algebra. Furthermore, $m(H_{\rho}^{0}\times H_{\rho})=H_{\rho}^{0}$.
@@
[48, Theorem 4.1] It holds that $H_{\rho^{\prime}}\subset H_{\rho}$ and the embedding operator
@@
[48, Lemma 4.3] It holds that $\mathcal{I}\in L(H_{\rho^{\prime}}^{0},H_{\rho}^{1})$ with $\|\mathcal{I}\|\leq C_{\rho,\rho^{\prime}}$, where
@@
We then consider the mapping $\mathcal{S}$ given by $\mathcal{S}h:=h\cdot\mathcal{I}h$.
@@
[22, Corollary 5.1.2] It holds that $\mathcal{S}\in{\rm Lip}^{{\rm loc}}(H_{\rho}^{0},H_{\rho}^{0})$ and there exists a constant $C>0$ such that $\|\mathcal{S}h\|_{\rho}\leq C\|h\|_{\rho}^{2}$, for all $h\in H_{\rho}^{0}$.
@@
Then, the product $\alpha:=\mathcal{S}\circ\beta=\beta\cdot\mathcal{I}\beta$ satisfies $\alpha\in{\rm Lip}^{{\rm loc}}(X,H_{\rho}^{0})\cap{\rm LG}(X,H_{\rho}^{0})$.
@@
and we denote by $W_{K}:\mathbb{R}_{+}\to\mathbb{R}_{+}$ its inverse. We also introduce the strictly increasing function
$w_{K}:\mathbb{R}_{+}\to\mathbb{R}_{+},\quad w_{K}(r):=W_{K}(r)\wedge r.$
@@
$B_{\Gamma}(r)$ $=B_{\mathcal{I}}(B_{\gamma}(r))=C_{\rho,\rho^{\prime}}B_{\gamma}(r).$
@@
Let $(E,\mathcal{E},F)$ be a measure space and $(Z,\mathcal{Z})$ a measurable space. Let $\gamma:X\times E\to H_{\rho^{\prime}}^{0}$ be a $\mathcal{B}(X)\otimes\mathcal{E}$-measurable function and $\lambda:Z\times E\to(0,\infty)$ a $\mathcal{Z}\otimes\mathcal{E}$-measurable function. Suppose that there exist a nonnegative function $\kappa\in L^{1}(F)\cap L^{2}(F)\cap L^{3}(F)$, an increasing function $M:\mathbb{R}_{+}\to\mathbb{R}_{+}$ and a function $\Lambda:Z\to\mathbb{R}_{+}$ such that the following conditions are satisfied:
@@
    (D.4) $B_{\gamma(\cdot,x)}(r)\leq w_{K_{\rho,\rho^{\prime}}}(\kappa(x)(1+r)),\qquad\text{ for all }r\in\mathbb{R}_{+};$
@@
$\int_{E}\|\gamma(h,x)\|_{\rho}^{2}F({\rm d}x)$ $\leq\int_{E}B_{\gamma(\cdot,x)}(r)^{2}F({\rm d}x)\leq\int_{E}w_{K_{\rho,\rho^{\prime}}}(\kappa(x)(1+r))^{2}F({\rm d}x)$
$\leq(1+r)^{2}\int_{E}\kappa(x)^{2}F({\rm d}x)\leq 2\int_{E}\kappa(x)^{2}F({\rm d}x)(1+\|h\|_{X}^{2}),$
@@
for all $r\in\mathbb{R}_{+}$. By (D.3), (D.4) and the inequalities $V_{K_{\rho,\rho^{\prime}}}(w_{K_{\rho,\rho^{\prime}}}(r))\leq r$ and $w_{K_{\rho,\rho^{\prime}}}(r)\leq r$, we obtain
$L_{\bar{\alpha}(z,\cdot,x)}(r)$ $\leq K\kappa(x)M(r)\big{(}1+\Lambda(z)\kappa(x)+\Lambda(z)\kappa(x)^{2}(1+r)\big{)},$
$B_{\bar{\alpha}(z,\cdot,x)}(r)$ $\leq K\big{(}\kappa(x)(1+r)+\Lambda(z)\kappa(x)^{2}(1+r)\big{)}=K\big{(}\kappa(x)+\Lambda(z)\kappa(x)^{2}\big{)}(1+r).$
@@
$\leq KM(r)\bigg{(}\int_{E}\kappa(x)F({\rm d}x)+\Lambda(z)\int_{E}\kappa(x)^{2}F({\rm d}x)+\Lambda(z)(1+r)\int_{E}\kappa(x)^{3}F({\rm d}x)\bigg{)}\|h-g\|_{X},$
@@
$\leq K\bigg{(}\int_{E}\kappa(x)F({\rm d}x)+\Lambda(z)\int_{E}\kappa(x)^{2}F({\rm d}x)\bigg{)}(1+\|h\|_{X}).$

### v3

@@ 4.4. Existence of affine realizations
Observe first that $\widehat{X}_{\cdot\wedge 1}=\widehat{X}^{1}\in\mathcal{X}_{1,[0,1]}\subseteq\mathcal{X}_{1}$. Proceeding by induction, suppose that $\widehat{X}_{\cdot\wedge n}\in\mathcal{X}_{1,[0,n]}$, for some $n\in\mathbb{N}$. The definition of $\widehat{X}$ implies that
@@
(A.2) $=\mathbf{1}_{\{t<n\}}(1+\widehat{X}_{t\wedge n})+\mathbf{1}_{\{t\geq n\}}\frac{1+\widehat{X}_{n}}{1+\widehat{X}^{n+1}_{n}}(1+\widehat{X}^{n+1}_{t}).$
As argued at the beginning of this proof, the sets $1+\mathcal{X}^{\mathrm{fin}}_{1,[0,n]}$, for $n\in\mathbb{N}$, satisfy the requirements of [44, Definition 1.1]. Therefore, by the arguments given in [44, pages 1358–1359], fork-convexity also holds for their closures $1+\mathcal{X}_{1,[0,n]}$, for all $n\in\mathbb{N}$. Since $\widehat{X}_{\cdot\wedge n}\in\mathcal{X}_{1,[0,n]}\subseteq\mathcal{X}_{1,[0,n+1]}$ and $\widehat{X}^{n+1}\in\mathcal{X}_{1,[0,n+1]}$, equation (A.2) and fork-convexity of $1+\mathcal{X}_{1,[0,n+1]}$ imply that $\widehat{X}_{\cdot\wedge(n+1)}\in\mathcal{X}_{1,[0,n+1]}\subseteq\mathcal{X}_{1}$. We have thus shown that $\widehat{X}_{\cdot\wedge n}\in\mathcal{X}_{1}$, for all $n\in\mathbb{N}$. Since $\widehat{X}$ is a semimartingale and $\widehat{X}_{\cdot\wedge n}=\mathbf{1}_{[\kern-1.22911pt[0,n]\kern-1.22911pt]}\cdot\widehat{X}$, for all $n\in\mathbb{N}$, the closedness of $\mathcal{X}_{1}$ in the semimartingale topology implies that $\widehat{X}\in\mathcal{X}_{1}$.
Conversely, if a process $D$ is a supermartingale deflator and $X\in\mathcal{X}_{1}$, then the supermartingale property of $D(1+X)$ together with the fact that $D_{0}\leq 1$ implies that $0\leq\mathbb{E}[D_{T}(1+X_{T})]\leq 1$, for all $T\in\mathbb{R}_{+}$. This implies that the set $D_{T}\mathcal{X}_{1}(T)$ is bounded in probability and, hence, NUPBR holds.
Finally, suppose that $D$ is a local martingale deflator and let $n\in\mathbb{N}$ and $A\in\mathcal{A}^{n}$. By definition, $DX^{A}$ is an $\mathbb{R}^{n}$-valued local martingale. For every $H\in L(X^{A})$ with $X:=H\cdot X^{A}\geq-1$, the process $DX$ is a local martingale (see, e.g., [28, Lemma 4.2]). Since every non-negative local martingale with integrable initial value is a supermartingale, this implies that $D(1+X)$ is a supermartingale, meaning that $D$ is a supermartingale deflator for the set $\bigcup_{n\geq 1}\mathcal{X}^{n}_{1}$. By Fatou’s lemma, the same property holds for the closure $\mathcal{X}_{1}$, thus proving that $D$ is a supermartingale deflator. ∎
@@ Appendix B Conditional expectation with respect to a Doléans measure
We recall the notion of conditional expectation with respect to a Doléans measure, as introduced in [36] (see also [37, Section III.3c]). Let $\mu$ be an integer-valued random measure on $\mathbb{R}_{+}\times E$ with compensator $\nu$. The positive Doléans measure $M_{\mu}$ on $(\Omega\times\mathbb{R}_{+}\times E,\mathcal{F}\otimes\mathcal{B}(\mathbb{R}_{+})\otimes\mathcal{B}(E))$ is defined by
@@
Suppose there exists a sequence $\{\tau_{n}\}_{n\in\mathbb{N}}$ of stopping times increasing a.s. to infinity as $n\to+\infty$ such that $\mathbb{E}[(|\varphi|\ast\mu)_{\tau_{n}}]<+\infty$, for all $n\in\mathbb{N}$. For each $n\in\mathbb{N}$, the function $H_{n}:\Omega\times\mathbb{R}_{+}\times E\to\{0,1\}$ defined by $H_{n}:=\mathbf{1}_{[\kern-1.22911pt[0,\tau_{n}]\kern-1.22911pt]\times E}$ is $\widetilde{\mathcal{P}}-$measurable and bounded. Therefore, property (B.1) implies that
@@
for all $n\in\mathbb{N}$, where in the last equality we made use of [37, Theorem II.1.8]. This shows that the process $M_{\mu}[|\varphi||\widetilde{\mathcal{P}}]\ast\nu$ is locally integrable. The converse implication can be shown in the same way. To prove the second part of the lemma, by localization we can assume that $\mathbb{E}[(|\varphi|\ast\mu)_{\infty}]<+\infty$. Let $\tau$ be an arbitrary stopping time. Similarly as above, letting the function $H$ be defined by $H:=\mathbf{1}_{[\kern-1.22911pt[0,\tau]\kern-1.22911pt]\times E}$,
@@
thus implying that the process $\varphi\ast\mu-M_{\mu}[\varphi|\widetilde{\mathcal{P}}]\ast\nu$ is a martingale. In view of [37, Theorem I.3.18], this suffices to deduce that $M_{\mu}[\varphi|\widetilde{\mathcal{P}}]\ast\nu$ is the compensator of $\varphi\ast\mu$. ∎
@@ Appendix D Properties of multi-dimensional Filipović spaces
    the translation semigroup $(\mathcal{S}_{t})_{t\geq 0}$ is a $C_{0}$-semigroup on $H_{\rho}$;
@@
[53, Lemma 4.2] The pair $(H_{\rho},m)$, where $m:H_{\rho}\times H_{\rho}\to H_{\rho}$ denotes the pointwise multiplication $m(h,g):=hg=h\cdot g$, is a commutative algebra. Furthermore, $m(H_{\rho}^{0}\times H_{\rho})=H_{\rho}^{0}$.
@@
[53, Theorem 4.1] It holds that $H_{\rho^{\prime}}\subset H_{\rho}$ and the embedding operator
@@
[53, Lemma 4.3] It holds that $\mathcal{I}\in L(H_{\rho^{\prime}}^{0},H_{\rho}^{1})$ with $\|\mathcal{I}\|\leq C_{\rho,\rho^{\prime}}$, where
@@
We then consider the mapping $\mathbf{S}$ given by $\mathbf{S}h:=h\cdot\mathcal{I}h$.
@@
[22, Corollary 5.1.2] It holds that $\mathbf{S}\in{\rm Lip}^{{\rm loc}}(H_{\rho}^{0},H_{\rho}^{0})$ and there exists a constant $C>0$ such that $\|\mathbf{S}h\|_{\rho}\leq C\|h\|_{\rho}^{2}$, for all $h\in H_{\rho}^{0}$.
@@
Then, the product $\alpha:=\mathbf{S}\circ\beta=\beta\cdot\mathcal{I}\beta$ satisfies $\alpha\in{\rm Lip}^{{\rm loc}}(X,H_{\rho}^{0})\cap{\rm LG}(X,H_{\rho}^{0})$.
@@
and we denote by $W_{K}:\mathbb{R}_{+}\to\mathbb{R}_{+}$ its inverse. Note that $V_{K}(r)\geq r$ for all $r\in\mathbb{R}_{+}$, which implies $W_{K}(r)\leq r$ for all $r\in\mathbb{R}_{+}$.
@@
$B_{\Gamma}(r)$ $=B_{\mathcal{I}}(B_{\gamma}(r))=C_{\rho,\rho^{\prime}}B_{\gamma}(r),$
@@
Let $(E,\mathcal{E},F)$ be a measure space and $(Z,\mathcal{Z})$ a measurable space. Let $\gamma:X\times E\to H_{\rho^{\prime}}^{0}$ be a $\mathcal{B}(X)\otimes\mathcal{E}$-measurable function and $\lambda:Z\times E\to(0,\infty)$ a $\mathcal{Z}\otimes\mathcal{E}$-measurable function. Suppose that there exist a nonnegative function $\kappa\in L^{1}(F)\cap L^{2}(F)\cap L^{3}(F)$, an increasing function $M:\mathbb{R}_{+}\to\mathbb{R}_{+}$, a constant $N\in\mathbb{R}_{+}$ and a function $\Lambda:Z\to\mathbb{R}_{+}$ such that the following conditions are satisfied:
@@
    (D.4) $B_{\gamma(\cdot,x)}(r)\leq W_{K_{\rho,\rho^{\prime}}}(N\kappa(x)(1+r)),\qquad\text{ for all }r\in\mathbb{R}_{+};$
@@
$\int_{E}\|\gamma(h,x)\|_{\rho}^{2}F({\rm d}x)$ $\leq\int_{E}B_{\gamma(\cdot,x)}(r)^{2}F({\rm d}x)\leq\int_{E}W_{K_{\rho,\rho^{\prime}}}(N\kappa(x)(1+r))^{2}F({\rm d}x)$
$\leq N^{2}(1+r)^{2}\int_{E}\kappa(x)^{2}F({\rm d}x)\leq 2N^{2}\int_{E}\kappa(x)^{2}F({\rm d}x)(1+\|h\|_{X}^{2}),$
@@
for all $r\in\mathbb{R}_{+}$. By (D.3), (D.4) and the relations $V_{K_{\rho,\rho^{\prime}}}(W_{K_{\rho,\rho^{\prime}}}(r))=r$ and $W_{K_{\rho,\rho^{\prime}}}(r)\leq r$, we obtain
$L_{\bar{\alpha}(z,\cdot,x)}(r)$ $\leq K\kappa(x)M(r)\big(1+\Lambda(z)\kappa(x)+N\Lambda(z)\kappa(x)^{2}(1+r)\big),$
$B_{\bar{\alpha}(z,\cdot,x)}(r)$ $\leq K\big(N\kappa(x)(1+r)+N\Lambda(z)\kappa(x)^{2}(1+r)\big)=KN\big(\kappa(x)+\Lambda(z)\kappa(x)^{2}\big)(1+r),$
@@
$\leq KM(r)\bigg(\int_{E}\kappa(x)F({\rm d}x)+\Lambda(z)\int_{E}\kappa(x)^{2}F({\rm d}x)+N\Lambda(z)(1+r)\int_{E}\kappa(x)^{3}F({\rm d}x)\bigg)\|h-g\|_{X},$
@@
$\leq KN\bigg(\int_{E}\kappa(x)F({\rm d}x)+\Lambda(z)\int_{E}\kappa(x)^{2}F({\rm d}x)\bigg)(1+\|h\|_{X}).$